# Effects of Geometric Modelling and Blood Rheology in Patient-Specific Arterial Blood Flow Simulations with Speed-Accuracy Trade-Off Analysis

### v3

@@ 2.2.2 Blood viscosity models
    The Newtonian model [28] is widely employed in blood flow simulations, primarily due to its linearity and the resulting computational efficiency. It is expected that for arteries of relatively larger size, Newtonian model predictions are negligibly different from the non-Newtonian ones. In the present study, blood viscosity is taken to be a constant, namely
@@
    The Carreau-Yasuda [29] is one of the most widely used Generalized Newtonian fluid (GNF) model. For viscoelastic fluids, the Carreau-Yasuda model is more realistic than the Power-law model since it shows a plateau at both low and high shear rates. Specifically, blood demonstrates near-constant viscosity at high shear rates, consistent with the assumptions of a Newtonian fluid under such conditions. The Carreau-Yasuda viscosity model is given as:
@@
    The recently developed viscosity model by Apostolides and Beris [30] is based on the Casson viscosity model. The parameters are functions of the constituents of blood such as hematocrit concentration $\phi$ and fibrinogen cf, whose values are adjusted using various flow-related data available in the literature. The apparent fluid viscosity is given by the following equation:
@@
    Figure 3: Blood viscosity models selected for simulating the blood flow in the descending aorta. Newtonian, Carreau Yasuda, and Apostolides and Beris models are shown.
@@
    At fully developed conditions, the local shear rate is zero at the center of a smooth circular channel. This local shear rate will make the apparent viscosity infinitely large, leading to an instability in the numerical solver. To ensure the stability of the solver at fully developed conditions, Miller and Morris [31] suggested that the shear rate magnitude is regularised as $\dot{\gamma}=\dot{\gamma}+\dot{\gamma}_{r}$, where the non-local shear rate $\dot{\gamma}_{r}<<\dot{\gamma}$ and given as
@@
    Here, $U_{max}$ is the maximum velocity attained in the Newtonian viscosity model, $\epsilon=0.01$ is the regularization parameter, and $L_{c}$ is the characteristic length. In the present study, it is chosen as the inlet diameter of the Level 1 artery. The parameters selected are appropriate for a normal human body temperature of 310 K and summarized in Table 3.
    Table 3: Parameters considered for the Apostolidis and Beris blood viscosity model[30]
@@
3 Flow distribution in a descending aorta
3.1 Grid Independence Test and Validation
@@
3D simulations of the flow distribution in the descending aorta in patient-specific (Level 1) and its simplified versions (Levels 2 and 3) are performed using the finite volume method within ANSYS-Fluent®, a commercial flow solver. For the grid independence test for the three geometries, the steady flow problem is solved on four refined meshes (M1, M2, M3, and M4). The solver used in this study is also validated with the results of Nagargoje et al. [32] for a symmetric bifurcation with pulsatile inlet boundary conditions. The pulsatile waveform and the geometry of interest are shown in Figure 4. In Figure 5(a), the average velocity obtained is shown as a function of the inverse of the mesh element size. Following grid independence, the M3 mesh is selected for further simulations. Validation against the work of [32] is shown in Figure 5(b). The $z$-component velocity is plotted on the centerline near the bifurcation for three-time instants of the oscillatory inflow condition. The validation results are seen to be satisfactory, which provides a check on the correctness of the solver.
@@
Figure 4: (a) Oscillatory velocity inlet profile. (b) symmetric bifurcation geometry used by Nagargoje et al. [32]
@@
Next, flow metrics and hemodynamic parameters are analyzed to understand the implications of the geometric simplifications for the three viscosity models. Quantities of interest are the velocity profiles and wall shear stress, along with the hemodynamic parameters such as the time-averaged wall shear stress (TAWSS) and Oscillatory Shear Index (OSI). Analysis in terms of the Q-criterion for strain rate and rotation are also reported.
@@
Figure 5: (a) Mesh independence test is performed for Levels 1-3 of the descending aorta. The average velocity at the outlet is presented as a function of the inverse of the mesh element size. Here, $e$ is the mesh element size, and a higher inverse element size indicates a finer mesh. (b) The $z$-velocity component (w) is plotted at time instants $t_{1},t_{2},$ and $t_{3}$ at the bifurcation of the geometry and matched against comparable results of Nagargoje et al. [32]
3.2 Velocity Profiles
@@
This section presents the cross-sectional velocity distribution during the systolic and diastolic phases over various planes of the descending aorta for the three viscosity models and Levels 1-3 geometries. The velocity data are shown in Figures 6 and 7 for the 4th cycle of simulation since the flow was shown to exhibit clear periodicity after the third cycle, indicating a fully developed dynamic steady state. Specific time instants from the cycle, as highlighted in left-most columns of Figures 6 and 7 are used for discussion.
@@
The systolic phase of the waveform (Figure 2) includes three key instants: $t_{1}$ = 0.2 s in the initial instant of increase in inflow velocity, $t_{2}$ = 0.3 s is the peak, and $t_{3}$= 0.45 s is the deceleration phase. The diastolic phase is analyzed at two time instants, $t_{4}$ = 0.6s and $t_{5}$ = 0.7s, representing flow fluctuation towards the end of the cycle. This section further describes the velocity characteristics over the cross-section in the form of centerline plots along with streamlines, contours, and vectors for the selected viscosity models and geometric simplification levels. In Figures 6 and 7, the distribution of the axial velocity component (denoted by $u$) is shown at the centerline of the cross-section located at a distance of $z=16D$ from the extended inlet as a function of the normalized $y$-coordinate. We show a comparison across two factors; in the first set of figures (Figure 6), we focus on comparing the viscosity models for a particular level of simplification, which shows subtleties in model predictions. In Figure 7, we show a comparison across the level of simplification for any given viscosity model. This completes the picture and shows how modeling aspects are coupled with geometry details. In Figure 6, each level of geometry simplification is shown in a column, while rows represent distinct phases of the pulsatile flow cycle. The velocity profiles, normalized across the channel width, are depicted by red, blue, and black lines for the Newtonian, Carreau-Yasuda, and Apostolidis-Beris models, respectively. The axial velocity component distribution is shown at the selected time instants for the individual viscosity models for a particular geometry simplification level. An asymmetric velocity distribution is observed for Level 1 only. This is expected since the perturbation and undulations across the circumference at any axial position are retained in Level 1, which leads to an asymmetric flow in succeeding cross-sections. Interestingly, in Level 1, the flow profiles are nearly identical for all phases of the cycle for the Newtonian and Carreau-Yasuda viscosity models. This similarity is expected since the Newtonian and Carreau-Yasuda models predict similar viscosities at high shear rates (Figure 3). The velocity predictions for the Apostolidis and Beris viscosity model are slightly larger than the other two for all the phases. For Level 2 (Figure 6(b)), the geometry is simplified and made symmetric around the circumference, and symmetric flow profiles are observed for all the viscosity models at every phase of the pulsatile flow. Interestingly, at this level, the two non-Newtonian models - Carreau-Yasuda and Apostolidis and Beris- predict similar velocity profiles, whereas the Newtonian predictions show an overshoot near the centerline. Note that such differences at the centerline are known between Newtonian and non-Newtonian models for flows in a smooth duct where the latter needs a resolution of the shear rate singularity within the definition of apparent viscosity. For the simplest Level 3 in Figure 6(c) which is a tapered smooth tube, the predictions across viscosity models are largely symmetric and similar to each other, owing to the symmetry of the geometry.
Next, in Figure 7, we show the velocity comparison across geometry simplifications for each viscosity model. The data shows an accuracy check for such simplifications across different viscosity models considered in the present study. Columns (a), (b), and (c) depict the velocity profiles for the Newtonian, Carreau-Yasuda, and Apostolidis and Beris viscosity models, respectively. As mentioned earlier, the rows represent distinct phases of the pulsatile flow cycle. Largely, the trends agree with each other, except for the asymmetry that is only visible for the patient-specific geometry (Level 1). The largest differences are observed at the start of the cycle (first row) and during the deceleration phase (third row). Interestingly, for the deceleration phase, only the predictions with the combination of Level 3 and Newtonian fluid are significantly more than the others. Thus, with respect to velocity profiles, the simplified geometry can capture the trends quite well for all viscosity models, except for the expected asymmetry in Level 1.
In what follows, further analyses in terms of streamlines, velocity contours, and vectors are provided. Figure 8a shows velocity magnitude contours at various cross-sectional planes in the descending aorta in the Level 1 geometry for various phases of the inflow waveform. Additionally, Figure 8b displays the velocity magnitude combined with streamline plots over the medial plane of the Level 1 aorta. These details are presented here for the Apostolidis and Beris viscosity model. The key trends remain similar to the other two models and hence, are not repeated. During the systolic phase, the inflow increases steadily, reaching a peak corresponding to the maximum flow rate. This progressive increase in inflow elevates wall forces, consequently enhancing wall loading. In contrast, the diastolic phase exhibits a much more complex flow behavior, driven by the adverse pressure gradient that develops as the inflow decelerates in time. The adverse gradient generates regions of backflow where the flow reverses direction near the walls.
@@
These regions are particularly pronounced in areas with irregular geometry or abrupt changes in cross-section and surface perturbations. The backflow generated disrupts the unidirectional flow and interacts with the surrounding fluid, leading to flow separation and the formation of recirculation zones. The interplay between backflow, flow circulation, and pressure gradients indicates the sensitivity of the flow field to boundary conditions and geometric variations and reinforces the importance of retaining details in the geometry. Thus, a prediction accuracy loss is expected with geometry simplifications, some of which are examined next in Figure 9.
Streamlines in the three geometries at all time instants are shown in Figure 9 with the Apostolidis and Beris viscosity model. It is observed that during the systolic phase, all the streamlines are unidirectional and no flow circulation is observed. However, as the flow enters the diastolic regime, flow circulations appear in Level 1 as well as Level 2 but are completely absent in Level 3. These flow recirculation zones are a result of the interplay of the geometrical perturbations and adverse pressure gradient conditions. With the absence of perturbations in Level 3, such zones of backflow are absent. Reversed flow zones are also larger in Level 1, relative to Level 2, owing to geometric variations coupled with a loss of symmetry. Thus, as the artery geometry is simplified, the intensity of circulation decreases and disappears in the limit of the smooth geometry approximation, namely, Level 3. Flow circulation patterns are important from a medical point of view since they affect the wall shear stress (WSS) and pressure loading. The circulation region being stagnant will promote particulate deposition leading to thrombosis. Hence, delineating such regions is important from a biomedical perspective.
a
b
@@
3.3 Hemodynamic Parameters
3.3.1 Wall shear stress (WSS)
@@
Wall shear stress (WSS) plays a crucial role in regulating the endothelial function and has been widely recognized as a key factor in the development of atherosclerosis. Endothelial cells sense the wall shear stress via multiple responsive cell components, and activation or deactivation of intra-cellular pathways leads to gene and protein expression regulation[33]. This continuous process is important for remodeling, development, and growth of the vascular walls. Moreover, clinical significance of wall shear stress lies in its ability to predict the risk of future cardiovascular events, making it an important biomarker in cardiovascular health. Mathematically, WSS can be calculated with the help of velocity gradients $(\nabla\textbf{u})$ as follows:
$\bar{\bar{\tau}}_{w}=2\mu_{app}\bar{\dot{\gamma}}$ (9)
@@
The strain rate ($\bar{\bar{\gamma}}$) is given by the sum of the velocity gradient and the transpose of the velocity gradient $\frac{1}{2}(\nabla\textbf{u}+\nabla\textbf{u}^{T})$ at the wall. The magnitude of the strain rate tensor field can be calculated as $\dot{\gamma}=\sqrt{2\bar{\bar{\gamma}}\colon\bar{\bar{\gamma}}}$. Wall shear stress is calculated for the patient-specific and simplified geometries (all levels) and presented in the form of contours. The wall shear stress (WSS) contours are shown in Figure 10 for all the time instants of the pulsatile cycle for the original and other geometric simplification levels while adopting the Apostolidis and Beris viscosity model. The features of WSS in Level 1 are particularly significant in predicting the onset and progression of disease, as they provide detailed insights, as well as the identification of the critical regions.
@@
Figure 10: Wall shear stress (WSS) contours superimposed with streamlines are shown at the selected time instants of the vascular cycle for the Apostolidis and Beris viscosity model. Within (i) to (v), Levels 1 to Level 3 are shown from the left.
@@
Figure 10 shows how the simulation accuracy deteriorates by simplifying the geometry. Firstly, for Level 1, with all perturbations preserved, we observe high variability in WSS at the peak instant of flow. At all times, the surface perturbations and the asymmetry generate several high (and low) stress zones of WSS. Some of these variabilities are captured in Level 2, although the asymmetry of Level 1 is lost. The high and low-stress zones appear as bands around the circumference, with an overall reduction of the variations observed in the WSS distributions of Level 1. This is most pronounced at the time instant of peak velocity. However, the WSS profile is smooth at Level 3 (tube of decreasing area), and all the variability is lost. Thus, Level 2 might still provide satisfactory predictions of WSS, which is entirely lost in Level 3. The variability in these hemodynamic parameters, in terms of the maximum and minimum values, are shown in Table 4 in a tabular form. For simplicity, for all three viscosity models, we show the data only at two time instants of the cycle - t = 0 s and t = 0.35 s (peak). The table further highlights the discussion about the accuracy of Levels 2 and 3 (relative to level 1) presented above. The table further shows that while the values in Level 2 retain some similarity, the values for Level 3 are significantly different in most cases. These trends broadly fit the discussion presented above.
@@
Table 4: Minimum and maximum wall shear stress magnitude values recorded at the start (t = 0 s) and at the peak (t = 0.35 s) of the cardiac cycle for the simplified geometrical levels and tabulated for the selected viscosity models.

### v4

@@ II.2.2 Blood viscosity models
    The Newtonian model [merrill1967viscosity] is widely employed in blood flow simulations, primarily due to its linearity and the resulting computational efficiency. It is expected that for arteries of relatively larger size, Newtonian model predictions are negligibly different from the non-Newtonian ones. In the present study, blood viscosity is taken to be a constant, namely
@@
    The Carreau-Yasuda [cho1991effects] is one of the most widely used Generalized Newtonian fluid (GNF) models. For viscoelastic fluids, the Carreau-Yasuda model is more realistic than the Power-law model, as it exhibits a plateau at both low and high shear rates. Specifically, blood demonstrates near-constant viscosity at high shear rates, consistent with the assumptions of a Newtonian fluid under such conditions. The Carreau-Yasuda viscosity model is given as:
@@
    The recently developed viscosity model by Apostolides and Beris [apostolidis2014modeling] is based on the Casson viscosity model. The parameters are functions of the constituents of blood, such as hematocrit concentration $\phi$ and fibrinogen cf, whose values are adjusted using various flow-related data available in the literature. The apparent fluid viscosity is given by the following equation:
@@
    Figure 3: Blood viscosity models selected for simulating the blood flow in the descending aorta. Newtonian, Carreau Yasuda, and Apostolides & Beris models are shown.
@@
    At fully developed conditions, the local shear rate is zero at the center of a smooth circular channel. This local shear rate will make the apparent viscosity infinitely large, leading to an instability in the numerical solver. To ensure the stability of the solver at fully developed conditions, Miller and Morris [miller2006normal] suggested that the shear rate magnitude is regularised as $\dot{\gamma}=\dot{\gamma}+\dot{\gamma}_{r}$, where the non-local shear rate $\dot{\gamma}_{r}<<\dot{\gamma}$ and given as
@@
    Here, $U_{max}$ is the maximum velocity attained in the Newtonian viscosity model, $\epsilon=0.01$ is the regularization parameter, and $L_{c}$ is the characteristic length. In the present study, it is chosen as the inlet diameter of the Level 1 artery is chosen. The parameters selected are appropriate for a normal human body temperature of 310 K and are summarized in Table 3.
    Table 3: Parameters considered for the Apostolidis and Beris blood viscosity model[apostolidis2014modeling]
@@
III Flow distribution in a descending aorta
III.1 Grid Independence Test and Validation
@@
3D simulations of the flow distribution in the descending aorta in patient-specific (Level 1) and its simplified versions (Levels 2 and 3) are performed using the finite volume method within ANSYS-Fluent®, a commercial flow solver. For the grid independence test for the three geometries, the steady flow problem is solved on four refined meshes (M1, M2, M3, and M4). The solver used in this study is also validated with the results of Nagargoje et al. [nagargoje2021pulsatile] for a symmetric bifurcation with pulsatile inlet boundary conditions. The pulsatile waveform and the geometry of interest are shown in Figure 4.
A steady flow problem is used for the grid independence test, instead of a pulsatile flow input. This is performed as a measure to save computational time and fix mesh sizes at the start, since pulsatile flow simulations incur much higher costs. The steady uniform inlet velocity (which itself is the average of the pulsatile waveform) is given to the physical domain of all arteries. Since the goal was to find the spatially converged grid, a single parameter is used to compare the results. The outlet average velocity is the simplest yet effective parameter used in various studies. In CFD simulations of incompressible flow, the SIMPLE algorithm (or its variant) does not exactly satisfy the mass balance constraint (i.e., the equation $\nabla\cdot u=0$ is not exactly solved). Hence, it is necessary to examine the divergence-free condition at the cell level (here 1E-5) and jointly check if the inflow and outflow values of average velocity match to within 10-3 % at all times. The former constraint at the cell level is also implemented in pulsatile flow simulations. In Figure 5(a), the average velocity obtained is shown as a function of the inverse of the mesh element size. Following grid independence, the M3 mesh is selected for further simulations. Validation against the work of [nagargoje2021pulsatile] is shown in Figure 5(b). The $z$-component velocity is plotted on the centerline near the bifurcation for three time instants of the oscillatory inflow condition. The validation results are satisfactory, providing a check on the solver’s correctness.
@@
Figure 4: (a) Oscillatory velocity inlet profile. (b) symmetric bifurcation geometry used by Nagargoje et al. [nagargoje2021pulsatile]
@@
Next, flow metrics and hemodynamic parameters are analyzed to understand the implications of the geometric simplifications for the three viscosity models. Quantities of interest include the velocity profiles and wall shear stress, along with hemodynamic parameters such as the time-averaged wall shear stress (TAWSS) and the Oscillatory Shear Index (OSI). Analysis in terms of the Q-criterion for strain rate and rotation is also reported.
@@
Figure 5: (a) Mesh independence test is performed for Levels 1-3 of the descending aorta. The average velocity at the outlet is presented as a function of the inverse of the mesh element size. Here, $e$ is the mesh element size, and a higher inverse element size indicates a finer mesh. (b) The $z$-velocity component (w) is plotted at time instants $t_{1},t_{2},$ and $t_{3}$ at the bifurcation of the geometry and matched against comparable results of Nagargoje et al. [nagargoje2021pulsatile]
III.2 Velocity Profiles
@@
This section presents the cross-sectional velocity distribution during the systolic and diastolic phases over various planes of the descending aorta for the three viscosity models and the Level 1-3 geometries. The velocity data are shown in Figures 6 and 7 for the 4th cycle of simulation, as the flow was observed to exhibit apparent periodicity after the third cycle, indicating a fully developed dynamic steady state. Specific time instants from the cycle, as highlighted in the left-most columns of Figures 6 and 7, are used for discussion.
@@
The systolic phase of the waveform (Figure 2) includes three key instants: $t_{1}$ = 0.2 s in the initial instant of increase in inflow velocity, $t_{2}$ = 0.3 s is the peak (of accelerating phase), and $t_{3}$= 0.45 s is the deceleration phase [owais2023effect]. The diastolic phase is analyzed at two time instants, $t_{4}=0.6$ s and $t_{5}=0.7$ s, representing flow fluctuations towards the end of the cycle. This section further describes the velocity characteristics across the cross-section in the form of centerline plots, along with streamlines, contours, and vectors, for the selected viscosity models and geometric simplification levels. In Figures 6 and 7, the distribution of the axial velocity component (denoted by $u$) is shown at the centerline of the cross-section located at a distance of $z=16D$ from the extended inlet as a function of the normalized $y$-coordinate. We present a comparison across two factors. In the first set of figures (Figure 6), we focus on comparing the viscosity models for a particular level of simplification, highlighting subtleties in model predictions. In Figure 7, we show a comparison across the level of simplification for any given viscosity model. This completes the picture, showing how modeling aspects are coupled with geometric details. In Figure 6, each level of geometry simplification is shown in a column, while rows represent distinct phases of the pulsatile flow cycle. The velocity profiles, normalized across the channel width, are depicted by red, blue, and black lines for the Newtonian, Carreau-Yasuda, and Apostolidis-Beris models, respectively. The axial velocity component distribution is shown at selected time instants for the individual viscosity models, corresponding to a particular level of geometry simplification. An asymmetric velocity distribution is observed for Level 1 only. This is expected since the perturbation and undulations across the circumference at any axial position are retained in Level 1, which leads to an asymmetric flow in succeeding cross-sections. Interestingly, in Level 1, the flow profiles are nearly identical for all phases of the cycle for the Newtonian and Carreau-Yasuda viscosity models. This similarity is expected since the Newtonian and Carreau-Yasuda models predict similar viscosities at high shear rates (Figure 3). The velocity predictions for the Apostolidis and Beris viscosity model are slightly larger than the other two for all the phases. For Level 2 (Figure 6(b)), the geometry is simplified and made symmetric around the circumference, and symmetric flow profiles are observed for all the viscosity models at every phase of the pulsatile flow. Interestingly, at this level, the two non-Newtonian models - Carreau-Yasuda and Apostolidis and Beris- predict similar velocity profiles, whereas the Newtonian predictions show an overshoot near the centerline. Note that such differences at the centerline are known between Newtonian and non-Newtonian models for flows in a smooth duct, where the latter needs a resolution of the shear rate singularity within the definition of apparent viscosity. For the simplest Level 3 in Figure 6(c), which is a tapered smooth tube, the predictions across viscosity models are largely symmetric and similar to each other, owing to the symmetry of the geometry. Next, in Figure 7, we present a comparison of velocities across different geometry simplifications for each viscosity model. The data shows an accuracy check for such simplifications across different viscosity models considered in the present study. Columns (a), (b), and (c) depict the velocity profiles for the Newtonian, Carreau-Yasuda, and Apostolidis and Beris viscosity models, respectively. As mentioned earlier, the rows represent distinct phases of the pulsatile flow cycle. Essentially, the trends agree with each other, except for the asymmetry that is only visible for the patient-specific geometry (Level 1). The most significant differences are observed at the start of the cycle (first row) and during the deceleration phase (third row). Interestingly, for the deceleration phase, only the predictions with the combination of Level 3 and Newtonian fluid are significantly more than the others. Thus, with respect to velocity profiles, the simplified geometry can capture the trends quite well for all viscosity models, except for the expected asymmetry in Level 1. The following sections present further analyses, including streamlines, velocity contours, and vectors. Figure 8(a) shows velocity magnitude contours at various cross-sectional planes in the descending aorta in the Level 1 geometry for multiple phases of the inflow waveform. Additionally, Figure 8(b) displays the velocity magnitude combined with streamline plots over the medial plane of the Level 1 aorta. These details are presented here for the Apostolidis and Beris viscosity model. The key trends remain similar to those of the other two models and are therefore not repeated. During the systolic phase, the inflow increases steadily, reaching a peak corresponding to the maximum flow rate. This progressive increase in inflow elevates wall forces, consequently enhancing wall loading. In contrast, the diastolic phase exhibits a much more complex flow behavior, driven by the adverse pressure gradient that develops as the inflow decelerates in time. The adverse gradient generates regions of backflow where the flow reverses direction near the walls.
@@
These regions are particularly pronounced in areas with irregular geometry, abrupt changes in cross-section, and surface perturbations. The backflow generated disrupts the unidirectional flow and interacts with the surrounding fluid, leading to flow separation and the formation of recirculation zones. The interplay between backflow, flow circulation, and pressure gradients highlights the sensitivity of the flow field to boundary conditions and geometric variations, reinforcing the importance of retaining detailed geometry. Thus, a loss in prediction accuracy is expected with geometry simplifications, some of which are examined next in Figure 9. Streamlines in the three geometries at all time instants are shown in Figure 9 with the Apostolidis and Beris viscosity model. It is observed that during the systolic phase, all the streamlines are unidirectional, and no flow circulation is observed. However, as the flow enters the diastolic regime, flow circulations appear in both Level 1 and Level 2, but are absent in Level 3. These flow recirculation zones result from the interplay of geometrical perturbations and adverse pressure gradient conditions. With the absence of perturbations in Level 3, such zones of backflow are absent. Reversed flow zones are also larger in Level 1, relative to Level 2, owing to geometric variations coupled with a loss of symmetry. Thus, as the artery geometry is simplified, the intensity of circulation decreases and disappears in the limit of the smooth geometry approximation, namely, Level 3. Flow circulation patterns are essential from a medical perspective, as they influence the wall shear stress (WSS) and pressure loading. A stagnant circulation region will promote particulate deposition, leading to thrombosis. Hence, delineating such areas is essential from a biomedical perspective. The pressure variation along different arterial levels is illustrated in Figure 10 for selected time instants of the cardiac cycle. The pressure values are presented relative to the physiological human blood-pressure range. It is observed that the geometric simplification of the descending aorta has a minimal influence on pressure, as the overall trend remains consistent at each time instant. However, the expected variation in pressure associated with changes in the inlet condition is clearly visible.
@@
Figure 10: Pressure contour shown for the simplified geometries for the Apostolidis and Beris viscosity model at the selected time instant.
III.3 Hemodynamic Parameters
III.3.1 Wall shear stress (WSS)
@@
Wall shear stress (WSS) plays a crucial role in regulating the endothelial function and has been widely recognized as a key factor in the development of atherosclerosis. Endothelial cells sense the wall shear stress via multiple responsive cell components, and activation or deactivation of intracellular pathways leads to gene and protein expression regulation[gijsen2019expert]. This continuous process is essential for remodeling, development, and growth of the vascular walls. Moreover, the clinical significance of wall shear stress lies in its ability to predict the risk of future cardiovascular events, making it an important biomarker in cardiovascular health. Mathematically, the shear stress can be calculated with the help of velocity gradients $(\nabla\textbf{u})$ as follows:
$\bar{\bar{\tau}}=2\mu_{app}\bar{\bar{\gamma}}$ (9)
The strain rate ($\bar{\bar{\gamma}}$) is given by the sum of the velocity gradient and the transpose of the velocity gradient $\bar{\bar{\gamma}}=(\nabla\textbf{u}+\nabla\textbf{u}^{T})/2$ at the wall.
For the wall shear stress, the traction vector of the shear stress is calculated by taking the wall unit normal projection as follows:
$\bar{t}=\bar{\bar{\tau}}\cdot\hat{n}$ (10)
Here, $\hat{n}$ is the unit normal vector of the artery wall.
The wall shear stress vector is calculated as the tangential components of the traction vector,
$\bar{\tau}_{w}=\bar{t}-(\bar{t}\cdot\hat{n})\hat{n}$ (11)
@@
The magnitude of the strain rate tensor field can be calculated as $\dot{\gamma}=\sqrt{2\bar{\bar{\gamma}}\colon\bar{\bar{\gamma}}}$. Wall shear stress is calculated for both patient-specific and simplified geometries (at all levels) and presented in the form of contours. The wall shear stress (WSS) contours are shown in Figure 11 for all the time instants of the pulsatile cycle for the original and other geometric simplification levels while adopting the Apostolidis and Beris viscosity model. The features of WSS in Level 1 are particularly significant in predicting the onset and progression of disease, as they provide detailed insights and identify critical regions.
@@
Figure 11: Wall shear stress (WSS) contours superimposed with streamlines are shown at the selected time instants of the vascular cycle for the Apostolidis and Beris viscosity model. Within (i) to (v), Levels 1 to 3 are shown from the left.
@@
Figure 11 illustrates how the simulation accuracy deteriorates as the geometry is simplified. Firstly, for Level 1, with all perturbations preserved, we observe high variability in WSS at the peak instant of flow. At all times, the surface perturbations and the asymmetry generate several high (and low) stress zones of WSS. Some of these variabilities are captured in Level 2, although the asymmetry of Level 1 is lost. The high and low-stress zones appear as bands around the circumference, with an overall reduction of the variations observed in the WSS distributions of Level 1. This is most pronounced at the time instant of peak velocity. However, the WSS profile is smooth at Level 3 (tube of decreasing area), and all the variability is lost. Thus, Level 2 might still provide satisfactory predictions of WSS, which is entirely lost in Level 3. The variability in these hemodynamic parameters, in terms of maximum and minimum values, is presented in Table 4. The calculated wall shear stress falls within a realistic range. Note that the time-averaged wall shear stress values fall within the limits of the study by Moore Jr. et al.[moore1994fluid]. For simplicity, we present the data for all three viscosity models at only two time instants of the cycle: t = 0 s and t = 0.35 s (peak). The table further highlights the discussion about the accuracy of Levels 2 and 3 (relative to Level 1) presented above. The table shows that while the values in Level 2 retain some similarity, the values for Level 3 are significantly different in most cases. These trends broadly fit the discussion presented above.
@@
Table 4: Minimum and maximum wall shear stress magnitude values (in Pa) recorded at the start (t = 0 s) and at the peak (t = 0.35 s) of the cardiac cycle for the simplified geometrical levels and tabulated for the selected viscosity models.